At what temperature is the speed of sound in helium (ideal gas, = 1.67, atomic mass = 4.003 u) the same as its speed in oxygen gas at 15.9 oC? The speed of sound in oxygen at 15.9°C is 322 m/s.
Answer #1

the speed of sound in helium V = sqrt(ykT/m)

where y = 1.67 , k = 1.38x10^-23 , V = 322 m/s

AND [​IMG]

so, T = (322^2 x 6.647x10^-27)/(1.37x1.38x10^-23) = 29.90 degree celsius = 30 oC approx ............ans